# Revisiting Adversarial Patch Defenses on Object Detectors: Unified Evaluation, Large-Scale Dataset, and New Insights

### v1

@@ 2.3 Adversarial Defenses on Object Detection
Objectseeker [47] (S&\&&P’23) Certified Defenses
@@ 3.1 Adversarial Patch Generation
In this paper, we generate adversarial patches, denoted as $\delta$, by conducting white-box adversarial attacks on various object detectors. Let =$\mathcal{F}=\{f_{1},f_{2},\ldots,f_{n}\}$ be the set of object detectors, where a single detector $f_{i}:X\longmapsto Y$ takes input images $x\in X$ and label $y\in Y$. We apply the adversarial patch $\delta$ to the input image $x$ through an applier $\mathcal{A}$ with transformations $t\in T$ (e.g., rotation, scaling, etc.). In this work, we focus on universal adversarial patches, which can be expressed as:
@@ 4.1 Evaluation Metrics
[…]
@@ 4.3 Patch Defenses in the Physical World
Adversarial patch attacks in the physical world can lead to serious security incidents. However, prior work has insufficiently explored defense performance in the physical world. Therefore, this section conducts experiments to investigate this issue. We select 30 types of patches and test them under different light intensities (1$\rightarrow$5, gradually increasing), at varying distances (3m/6m/9m), and angles (-90° to 90°). Using an iPhone16pro, we capture 540 frames of adversarial images. We test each defense and record the results, as shown in Fig. 3. Our experiments show that attack effectiveness declines with distance (e.g., AP rises from 60% to 85% as distance increases from 0m to 9m). Defenses perform worst in the [-30°, 30°) range, with AP scores 30-50% lower than other angles. Attack effectiveness decreases while defense performance improves with increasing light intensity (levels 1-5). As shown in Fig. 4, without defenses, the patch can successfully attack the victim detector under different angles, distances, and background environments. When using LGS [31] or NAPGuard [44], the patch is effectively eliminated. Therefore, defenses that perform well in the digital domain also generally demonstrate robustness in the physical world. Surprisingly, PAD [16], which performs well in the digital domain, is surpassed by many other defense methods in the physical world. Observing the patch masks generated by the defense methods, we find that although PAD can detect and remove attack patches, it is prone to misclassifying humans as patches, leading to incorrect predictions by the detector.

### v2

@@ 2.3 Adversarial Defenses on Object Detection
Objectseeker [47] (S&\&P’23) Certified Defenses
@@ 3.1 Adversarial Patch Generation
In this paper, we generate adversarial patches, denoted as $\delta$, by conducting white-box adversarial attacks on various object detectors. Let $\mathcal{F}=\{f_{1},f_{2},\ldots,f_{n}\}$ be the set of object detectors, where a single detector $f_{i}:X\longmapsto Y$ takes input images $x\in X$ and label $y\in Y$. We apply the adversarial patch $\delta$ to the input image $x$ through an applier $\mathcal{A}$ with transformations $t\in T$ (e.g., rotation, scaling, etc.). In this work, we focus on universal adversarial patches, which can be expressed as:
@@ 4.1 Evaluation Metrics
$\text{SmIoU}=$ $\frac{\sum_{n=1}^{N}\text{TP}_{n}}{\sum_{n=1}^{N}(\text{FP}_{n}+\text{TP}_{n}+\text{FN}_{n})}$ (2)
$\text{NmIoU}=$ $\frac{1}{N}\sum_{n=1}^{N}\frac{\text{TP}_{n}}{\text{FP}_{n}+\text{TP}_{n}+\text{FN}_{n}}$ (3)
@@ 4.3 Patch Defenses in the Physical World
[…]